OpenStudy (erniellyd28):

add the polynomials (6x to the second+3x+2)+(9x to the second+9x)

TheSmartOne (thesmartone):

Remember, you can only add like terms what is 6x^2 + 9x^2 = ? 3x + 9x = ? Make it easier for you to see how to add the like terms: 6x^2 + 9x^2 = x^2(6 + 9) = ?? 3x + 9x = x(3 + 9) = ??

OpenStudy (erniellyd28):

(6+9)=15 (3+9)=12

TheSmartOne (thesmartone):

6x^2 + 9x^2 = x^2(6 + 9) = 15x^2 3x + 9x = x(3 + 9) = 12x so now we have 6x^2+3x+2)+(9x^2+9x) when we rearrange the terms 6x^2 + 9x^2 + 3x + 9x + 2 And you told me that 6x^2 + 9x^2 = 15x^2 so we get 15x^2 + 3x + 9x + 2 and that 3x + 9x = 12x , soo 15x^2 + 12x + 2 Ta-da. :)

OpenStudy (erniellyd28):

and thats my answer?

TheSmartOne (thesmartone):

yes :)